# Particle Semi-Implicit Variational Inference

### v1

@@ 4 Theoretical analysis
We are interested in the behaviour of the PVI flow (10). However, the key issue in its study is that the drift in PVI flow might lack the necessary continuity properties to analyze using the existing theory. In this section, we instead analyze the related gradient flow of the more regular functional
@@
for $\gamma>0$. A similar modified flow was also explored in Crucinio et al. (2024) for similar reasons; they found empirically that, at least when using a tamed Euler scheme, setting $\gamma=0$ did not cause problems in practice. Similarly, our experimental results for PVI found $\gamma=0$ did not have issues. To provide an additional measure of confidence in the reasonableness of this regularization and of the use of this functional as a proxy for $\mathcal{E}_{\lambda}$ we establish that the minima of $\mathcal{E}^{\gamma}_{\lambda}$ converge to those of $\mathcal{E}_{\lambda}$ in the $\lambda\rightarrow 0$ limit.
@@
The proof uses techniques from $\Gamma$-convergence theory (introduced by De Giorgi, see e.g. De Giorgi and Franzoni (1975); see Dal Maso (2012) for a good modern introduction) and can be found in Section F.1. The gradient flow of $\mathcal{E}^{\gamma}_{\lambda}$, which we term $\gamma$-PVI, is given by
@@
where $r^{\gamma}_{t}=\mathrm{Law}(Z^{\gamma}_{t})$ and $b^{\gamma}(\theta,r,\cdot)=-\nabla_{z}\delta_{r}\mathcal{E}^{\gamma}[\theta,r]+\lambda_{r}\nabla_{z}\log p_{0}$. The derivation follows similarly from that in Section 3.2, and is omitted for brevity. The use of $\gamma>0$ is crucial for establishing key regularity conditions in our analysis. We proceed by stating our assumptions.
@@
We assume that the kernel $k$ and its gradient is bounded and has $K_{k}$-Lipschitz gradient; i.e., there exist constants $B_{k},K_{k}\in\mathbb{R}_{>0}$ such that $|k{{}_{\theta}}(x|z)|+\|\nabla_{(\theta,x,z)}k{{}_{\theta}}(x|z)\|\leq B_{k}$, and $\|\nabla_{x}k{{}_{\theta}}(x|z)-\nabla_{x}k{{}_{\theta^{\prime}}}(x^{\prime}|z^{\prime})\|\leq K_{k}(\|(\theta,x,z)-(\theta^{\prime},x^{\prime},z^{\prime})\|)$ hold for all $\theta,\theta^{\prime}\in\Theta$, $z,z^{\prime}\in\mathbb{R}^{d_{z}}$, and $x,x^{\prime}\in\mathbb{R}^{d_{x}}$,
@@
We assume that $\phi$ has $K_{\phi}$-Lipschitz gradient and bounded gradient. In order words, there is $a_{\phi},b_{\phi},:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}_{>0}$ such that $\phi$ satisfies $\|\nabla_{(\theta,z)}\phi(z,\epsilon)-\nabla_{(\theta,z)}\phi_{\theta^{\prime}}(z^{\prime},\epsilon)\|_{F}\leq(a_{\phi}\|\epsilon\|+b_{\phi})(\|(\theta,z)-(\theta^{\prime},z^{\prime})\|)$ and $\|\nabla_{(\theta,z)}\phi_{\theta}(z,\epsilon)\|_{F}\leq(a_{\phi}\|\epsilon\|+b_{\phi})$ for all $(\theta,z),(\theta^{\prime},z^{\prime})\in\Theta\times\mathbb{R}^{d_{z}},\epsilon\in\mathbb{R}^{d_{x}}$, where $\|\cdot\|_{F}$ denotes the Frobenius norm. We also assume that $p_{k}$ has finite second moments.
Assumptions 2 and 3 are intimately connected; under some regularity conditions, one may imply the other but we shall abstain from this digression for the sake of clarity. They are quite mild and hold for popular kernels such as $k{{}_{\theta}}(x|z)=\mathcal{N}(x;\mu_{\theta}(z),\Sigma)$ under some regularity assumptions on $\mu$ and $\Sigma$ (which we show in App. C). These assumptions are key to establishing that the drift in Eq. 17 is Lipschitz continuous Prop. 11 (in the Appendix), from which, we establish the existence and uniqueness of the solutions of Eq. 17.
@@
Under Assumptions 1, 2 and 3 and assume that $\gamma>0$ and $\mathbb{E}_{p_{k}(\epsilon)}\|s^{\gamma}_{\theta,r}(z,\epsilon)-s_{p}(z,\epsilon)\|$ is bounded (where $s^{\gamma}_{\theta,r}:(z,\epsilon)\mapsto\nabla\log(q_{\theta,r}\circ\phi_{\theta}(z,\epsilon)+\gamma)$); then, given $(\theta_{0},r_{0})\in\mathcal{M}$ then the solutions to Eq. 17 exists and is unique.
@@
where ${(r^{\gamma}_{t})}^{\otimes M}=\prod_{i=1}^{M}(r^{\gamma}_{t})$; $q^{\gamma,M}_{t}=\mathrm{Law}(\{Z^{\gamma,M}_{t,m}\}_{m=1}^{M})$; and $\theta_{t}^{\gamma,M}$ and $Z^{\gamma,M}_{t,m}$ are solutions to
@@
The proof can be found in Section F.3. Having established desirable properties about $\gamma$-PVI flow in the form of existence and uniqueness as well as asymptotic justification for using particles, we now provide a numerical evaluation to demonstrate the efficacy of our proposal.
@@ 5 Experiments
In this section, we compare PVI against other semi-implicit VI methods. As described in the App. A, these include unbiased semi-implicit variational inference (UVI) of Titsias and Ruiz (2019), semi-implicit variational inference (SVI) of Yin and Zhou (2018), and the score matching approach (SM) of Yu and Zhang (2023). Through experiments, we show the benefits of optimizing the mixing distribution; we compare the effectiveness of PVI against other SIVI methods on a density estimation problem on toy examples and a high-dimensional Bayesian neural network regression problem. The details for reproducing experiments as well as computation information can be found in Table 3.
@@ 5.1 Impact of the mixing distribution
Figure 1: Comparison of PVI and PVIZero on a bimodal mixture of Gaussians for various kernels. The plot shows the density $q_{\theta,r}$ from PVI and PVIZero as well as the KDE plot of $100$ particles from PVI.
From Prop. 1, it can be said that ultimately current SIVI methods utilize (directly or indirectly) a fixed mixing distribution whilst PVI does not. We are interested in establishing whether there is any benefit to optimizing the mixing distribution. Intuitively, the mixing distribution can be utilized to express complex properties, such as multimodality, which the neural network kernel $k_{\theta}$ can then exploit. If the mixing distribution is fixed, this means that the neural network must learn to express these complex properties directly—which can be difficult (Salmona et al., 2022). This intuition turns out to hold, but for the kernel to be able to exploit an expressive mixing distribution, it must be designed well. To illustrate this, consider the distributions $\frac{1}{2}\mathcal{N}(\mu,I)+\frac{1}{2}\mathcal{N}(-\mu,I)$ for $\mu=\{1,2,4\}$ and following kernels: the “Constant” kernel $\mathcal{N}(z,I_{2})$; “Push” kernel $\mathcal{N}(f_{\theta}(z),\sigma_{\theta}^{2}I_{2})$; “Skip” kernel $\mathcal{N}(z+f_{\theta}(z),\sigma_{\theta}^{2}I_{2})$; and “LSkip” kernel $\mathcal{N}(Wz+f_{\theta}(z),\sigma_{\theta}^{2}I_{2})$ where $W\in\mathbb{R}^{2\times 2}$;. We compare the results from PVI and PVIZero (PVI with $h_{r}=0$ to result in a fixed $r\approx\mathcal{N}(0,I_{2})$) to emulate PVI with a fixed mixing distribution. As $\mu$ gets larger, the complexity of the kernel (or the mixing distribution) must grow to express this (e.g., see (Salmona et al., 2022, Corollary 2)).
Fig. 1 shows the resulting densities and the learnt mixing distribution of PVI and PVIZero for different kernels and various $\mu$. For the constant kernel, PVI can solve this problem by learning a complex mixing distribution to express the multimodality. However, for the push kernel, it can be seen that as $\mu$ gets larger PVI and PVIZero suffer from mode collapse which we suspect is due to the mode-seeking behaviour of using reverse KL and why prior SIVI methods utilized annealing methods (see Yu and Zhang (2023, Section 4.1)). As a remedy, we utilize a skip kernel which can be seen to improve both PVI and PVIZero. In particular, both PVI and PVIZero were able to successfully express the bimodality in $mu=2$; however, PVIZero falls short when $\mu=4$ while PVI can express the multimodality by learning a bimodal mixing distribution. Since Skip requires $d_{z}=d_{x}$, we show that LSkip (which removes the requirement) exhibits a similar behaviour to Skip.
@@ 5.2 Density estimation
Figure 2: Contour plots of the densities $q_{\theta,r}$ (in blue) against the true densities (in black) for various toy examples. We also plot the absolute difference in the density of $q_{\theta,r}$ and the true density, i.e., $|q_{\theta,r}-p|$.
We follow prior works (e.g., Yin and Zhou (2018)) and consider three toy examples whose densities are shown in Fig. 2 (which are written explicitly in Section H.2). In this setting, we use the kernel $k_{\theta}(x|z)=\mathcal{N}(x;z+\mu_{\theta}(z),\sigma_{\theta}^{2}I)$ with $d_{z}=d_{x}=2$ where $\mu_{\theta}(z)$ is a neural network whose architecture can be found in Section H.2. As a qualitative measure of performance, Fig. 2 shows the resulting approximating distribution of PVI which can be seen to be a close match to the desired distribution. To compare our proposal against prior work, we report the (sliced) Wasserstein distance (computed by POT (Flamary et al., 2021)) and the rejection power of a state-of-the-art two-sample kernel test (Biggs et al., 2023) between the approximating and true distribution in Table 1. The results reported are the average and standard deviation (from ten independent trials of the respective SIVI algorithms). In each trial, the rejection rate $p$ is computed from $100$ tests and the sliced Wasserstein distance is computed from $10000$ samples with $100$ projections. If the variational approximation matches the distribution, the rejection rate will be at the nominal level of $0.05$. It can be seen that PVI consistently obtains better performances than SIVI across all problems. PVI can achieve a rejection rate near nominal levels across all problems whilst other algorithms can achieve good performances on one but not the other. The details regarding how the Wasserstein distance is calculated and the hyperparameters used can be found in Section H.2.
@@
Banana $\bm{0.06}_{0.02}/{0.17}_{0.01}$ $\bm{0.05}_{0.02}/\bm{0.09}_{0.02}$ $0.13_{0.05}/0.31_{0.02}$ $0.59_{0.26}/0.29_{0.06}$
Multimodal $\bm{0.05}_{0.01}/\bm{0.05}_{0.01}$ $0.75_{0.18}/0.16_{0.07}$ $\bm{0.07}_{0.04}/0.07_{0.01}$ $0.59_{0.28}/0.16_{0.05}$
X-Shape $\bm{0.07}_{0.04}/\bm{0.07}_{0.02}$ $0.45_{0.32}/0.16_{0.07}$ $0.13_{0.05}/0.12_{0.01}$ $0.34_{0.20}/0.15_{0.05}$
Table 1: This table shows the rejection rate $p$ and (sliced) Wasserstein distance $w$ following the format $p/w$ (lower is better) with the subscripts showing the standard deviation estimated from $10$ independent runs. We indicate in bold when the rejection rate minus the standard deviation is lower than the nominal level $0.05$, and when the algorithm achieves the lowest Wasserstein score.

### v3

@@ 4 Theoretical analysis
We are interested in the behaviour of the PVI flow (10). However, a key issue in its study is that the drift in PVI flow might lack the necessary continuity properties to analyze using the existing theory. In this section, we instead analyze the related gradient flow of the more regular functional
@@
for $\gamma>0$. A similar modified flow was also explored in Crucinio et al., (2024) for similar reasons; they found empirically that, at least when using a tamed Euler scheme, setting $\gamma=0$ did not cause problems in practice. Similarly, our experimental results for PVI found $\gamma=0$ did not have issues. To provide an additional measure of confidence in the reasonableness of this regularization and of the use of this functional as a proxy for $\mathcal{E}_{\lambda}$, we establish that the minima of $\mathcal{E}^{\gamma}_{\lambda}$ converge to those of $\mathcal{E}_{\lambda}$ in the $\gamma\rightarrow 0$ limit.
@@
The proof uses techniques from $\Gamma$-convergence theory (introduced by De Giorgi, see e.g. De Giorgi and Franzoni, (1975); see Dal Maso, (2012) for a good modern introduction) and can be found in Section F.1. The gradient flow of $\mathcal{E}^{\gamma}_{\lambda}$, which we term $\gamma$-PVI, is given by
@@
where $r^{\gamma}_{t}=\mathrm{Law}(Z^{\gamma}_{t})$ and $b^{\gamma}(\theta,r,\cdot)=-\nabla_{z}\delta_{r}\mathcal{E}^{\gamma}[\theta,r]+\lambda_{r}\nabla_{z}\log p_{0}$. The derivation follows similarly to that in Section 3.2, and is omitted for brevity. The use of $\gamma>0$ is crucial for establishing key regularity conditions in our analysis. We proceed by stating our assumptions.
@@
We assume that the kernel $k$ and its gradient is bounded and has $K_{k}$-Lipschitz gradient; i.e., there exist constants $B_{k},K_{k}\in\mathbb{R}_{>0}$ such that $|k{{}_{\theta}}(x|z)|+\|\nabla_{(\theta,x,z)}k{{}_{\theta}}(x|z)\|\leq B_{k}$, and $\|\nabla_{x}k{{}_{\theta}}(x|z)-\nabla_{x}k{{}_{\theta^{\prime}}}(x^{\prime}|z^{\prime})\|\leq K_{k}(\|(\theta,x,z)-(\theta^{\prime},x^{\prime},z^{\prime})\|)$ hold for all $\theta,\theta^{\prime}\in\Theta$, $z,z^{\prime}\in\mathbb{R}^{d_{z}}$, and $x,x^{\prime}\in\mathbb{R}^{d_{x}}$.
@@
We assume that $\phi$ has Lipschitz gradient and bounded gradient. In other words, there is $a_{\phi}\in\mathbb{R}_{\geq 0},b_{\phi}\in\mathbb{R}_{>0}$ such that $\phi$ satisfies $\|\nabla_{(\theta,z)}\phi(z,\epsilon)-\nabla_{(\theta,z)}\phi_{\theta^{\prime}}(z^{\prime},\epsilon)\|_{F}\leq(a_{\phi}\|\epsilon\|+b_{\phi})(\|(\theta,z)-(\theta^{\prime},z^{\prime})\|)$ and $\|\nabla_{(\theta,z)}\phi_{\theta}(z,\epsilon)\|_{F}\leq(a_{\phi}\|\epsilon\|+b_{\phi})$ for all $(\theta,z),(\theta^{\prime},z^{\prime})\in\Theta\times\mathbb{R}^{d_{z}},\epsilon\in\mathbb{R}^{d_{x}}$, where $\|\cdot\|_{F}$ denotes the Frobenius norm. We also assume that $p_{k}$ has finite second moments.
Assumptions 2 and 3 are intimately connected; under some regularity conditions, one may imply the other but we shall abstain from this digression for the sake of clarity. Assumptions 2 and 3 are quite mild and hold for popular kernels such as $k{{}_{\theta}}(x|z)=\mathcal{N}(x;\mu_{\theta}(z),\Sigma)$ under some regularity assumptions on $\mu_{\theta}$ and $\Sigma$ (which we show in App. C). These assumptions are key to establishing that the drift in Eq. 19 is Lipschitz continuous (see Prop. 11 in the App.), from which, we establish the existence and uniqueness of the solutions of Eq. 19.
@@
Under Assumptions 1, 2 and 3, if $\gamma>0$ and $\mathbb{E}_{p_{k}(\epsilon)}\|s^{\gamma}_{\theta,r}(z,\epsilon)-s_{p}(z,\epsilon)\|$ is bounded (with $s^{\gamma}_{\theta,r}:(z,\epsilon)\mapsto\nabla_{x}\log(q_{\theta,r}\circ\phi_{\theta}(z,\epsilon)+\gamma)$); then, given $(\theta_{0},r_{0})\in\mathcal{M}$, the solutions to Eq. 19 exists and is unique.
@@
where ${(r^{\gamma}_{t})}^{\otimes M}=\prod_{i=1}^{M}(r^{\gamma}_{t})$; $q^{\gamma,M}_{t}=\mathrm{Law}(\{Z^{\gamma,M}_{t,m}\}_{m=1}^{M})$; $\theta_{t}^{\gamma,M}$ and $Z^{\gamma,M}_{t,m}$ are solutions to
@@
The proof can be found in Section F.3. Having established the existence and uniqueness of the $\gamma$-PVI flow, as well as an asymptotic justification for using particles, we now provide a numerical evaluation to demonstrate the efficacy of our proposal.
@@ 5 Experiments
In this section, we compare PVI against other semi-implicit VI methods. As described in the App. A, these include unbiased semi-implicit variational inference (UVI) of Titsias and Ruiz, (2019), semi-implicit variational inference (SVI) of Yin and Zhou, (2018), and the score matching approach (SM) of Yu and Zhang, (2023). Through experiments, we show the benefits of optimizing the mixing distribution; we compare the effectiveness of PVI against other SIVI methods on a density estimation problem on toy examples; and, we compare against other SIVI methods on posterior estimation tasks for (Bayesian) logistic regression and (Bayesian) neural network. The details for reproducing experiments as well as computation information can be found in Table 3. The code is available at https://github.com/jenninglim/pvi.
@@ 5.1 Impact of the mixing distribution
Figure 1: Comparison of PVI and PVIZero on a bimodal mixture of Gaussians for various kernels. The plot shows the density $q_{\theta,r}$ from PVI and PVIZero as well as the KDE plot of $r$ from PVI described by $100$ particles.
From Prop. 1, it can be said that ultimately current SIVI methods utilize (directly or indirectly) a fixed mixing distribution whilst PVI does not. We are interested in establishing whether there is any benefit to optimizing the mixing distribution. Intuitively, the mixing distribution can be utilized to express complex properties, such as multimodality, which the neural network kernel $k_{\theta}$ can then exploit. If the mixing distribution is fixed, this means that the neural network must learn to express these complex properties directly—which can be difficult (Salmona et al.,, 2022). This intuition turns out to hold, but for the kernel to exploit an expressive mixing distribution, it must be designed well. We illustrate this using the distributions $\frac{1}{2}\mathcal{N}(\mu,I)+\frac{1}{2}\mathcal{N}(-\mu,I)$ for $\mu=\{1,2,4\}$ and following kernels: the “Constant” kernel $\mathcal{N}(z,I_{2})$; “Push” kernel $\mathcal{N}(f_{\theta}(z),\sigma_{\theta}^{2}I_{2})$; “Skip” kernel $\mathcal{N}(z+f_{\theta}(z),\sigma_{\theta}^{2}I_{2})$; and “LSkip” kernel $\mathcal{N}(Wz+f_{\theta}(z),\sigma_{\theta}^{2}I_{2})$ where $W\in\mathbb{R}^{2\times 2}$;. We compare the results from PVI and PVIZero (PVI with $h_{r}=0$ to result in a fixed $r\approx\mathcal{N}(0,I_{2})$) to emulate PVI with a fixed mixing distribution. As $\mu$ gets larger, the complexity of the kernel (or the mixing distribution) must grow to express this (e.g., see (Salmona et al.,, 2022, Corollary 2)).
Fig. 1 shows the resulting densities and the learnt mixing distribution of PVI and PVIZero for different kernels and various $\mu$. For the constant kernel, PVI can solve this problem by learning a complex mixing distribution to express the multimodality. However, for the push kernel, it can be seen that as $\mu$ gets larger PVI and PVIZero suffer from mode collapse which we suspect is due to the mode-seeking behaviour of using reverse KL and why prior SIVI methods utilized annealing methods (see Yu and Zhang, (2023, Section 4.1)). As a remedy, we utilize a Skip kernel which can be seen to improve both PVI and PVIZero. In particular, both PVI and PVIZero were able to successfully express the bimodality in $\mu=2$; however, PVIZero falls short when $\mu=4$ while PVI can express the multimodality by learning a bimodal mixing distribution. Since Skip requires $d_{z}=d_{x}$, we show that LSkip (which removes the requirement) exhibits a similar behaviour to Skip.
@@ 5.2 Density estimation
Figure 2: Contour plots of the densities $q_{\theta,r}$ (in blue) against the true densities (in black) for various toy density estimation problems. We also plot the absolute difference in the density of $q_{\theta,r}$ and the true density, i.e., $|q_{\theta,r}-p|$.
We follow prior works (e.g., Yin and Zhou, (2018)) and consider three toy examples whose densities are shown in Fig. 2 (they are given explicitly in Section H.2). In this setting, we use the kernel $k_{\theta}(x|z)=\mathcal{N}(x;z+f_{\theta}(z),\sigma_{\theta}^{2}I)$ with $d_{z}=d_{x}=2$ where $f_{\theta}(z)$ is a neural network whose architecture can be found in Section H.2. As a qualitative measure of performance, Fig. 2 shows the resulting approximating distribution of PVI which can be seen to be a close match to the desired distribution. To compare methods quantitatively, we report the (sliced) Wasserstein distance (computed by POT (Flamary et al.,, 2021)) and the rejection power of a state-of-the-art two-sample kernel test (Biggs et al.,, 2023) between the approximating and true distribution in Table 1. The results reported are the average and standard deviation (from ten independent trials of the respective SIVI algorithms). In each trial, the rejection rate $p$ is computed from $100$ tests and the sliced Wasserstein distance is computed from $10000$ samples with $100$ projections. If the variational approximation matches the distribution, the rejection rate will be at the nominal level of $0.05$. It can be seen that PVI consistently performs better than SIVI across all problems. PVI can achieve a rejection rate near nominal levels across all problems whilst other algorithms can achieve good performances on one but not the other. The details regarding how the Wasserstein distance is calculated and the hyperparameters used can be found in Section H.2.
@@
Banana $\bm{0.06}_{0.02}/{0.17}_{0.01}$ $\bm{0.07}_{0.02}/\bm{0.11}_{0.03}$ $0.13_{0.05}/0.31_{0.02}$ $0.39_{0.24}/0.24_{0.12}$
Multimodal $\bm{0.05}_{0.01}/\bm{0.05}_{0.01}$ $0.65_{0.23}/0.16_{0.07}$ $0.13_{0.06}/0.08_{0.02}$ $0.14_{0.05}/0.10_{0.02}$
X-Shape $\bm{0.06}_{0.03}/\bm{0.07}_{0.01}$ $0.23_{0.16}/0.10_{0.04}$ $0.11_{0.04}/0.12_{0.01}$ $0.15_{0.11}/0.11_{0.03}$
Table 1: This table shows the rejection rate $p$ and average (sliced) Wasserstein distance $w$ for toy density estimation problems. It is written in the format $p/w$ (lower is better) with the subscripts showing the standard deviation estimated from $10$ independent runs. We indicate in bold when the rejection rate minus the standard deviation is lower than the nominal level $0.05$, and the algorithm that achieves the lowest Wasserstein score.
5.3 Bayesian logistic regression
As with others (Yin and Zhou,, 2018), we consider a Bayesian logistic regression problem on the waveform dataset (Breiman and Stone,, 1984). The model is expressed as $y\,|\,x,\bm{w}\sim\mathrm{Bernoulli}(\mathrm{Sigmoid}(\langle x,\overline{\bm{w}}\rangle))$ with prior $x\sim\mathcal{N}(0,0.01^{-1}\times I_{22})$ where $(y,\bm{w})\in\{0,1\}\times\mathbb{R}^{21}$ is the response and covariates, and $\overline{\bm{w}}:=[1,\bm{w}]$ is the covariates with appended one for the intercept. The “ground truth” is composed of posterior samples generated from running Markov chain Monte Carlo (MCMC) samples in Yin and Zhou, (2018). We use the kernel $k_{\theta}(x|z)=\mathcal{N}(Wz+f_{\theta}(z),\exp(\frac{1}{2}[M_{\theta}+M_{\theta}^{\top}]))$ where $\exp$ denotes the matrix exponential which ensures positive definiteness. In Fig. 3, we visually compare certain statistics of the distribution obtained from MCMC and distribution obtained from SIVI methods. Fig. 3a shows the pair-wise marginal posterior distributions for three weights $x_{1},x_{2},x_{3}$ chosen at random from MCMC and SIVI approximations; and in Fig. 3b we compare the correlation coefficients obtained from MCMC against SIVI methods. It can be seen that PVI obtains an approximation close to MCMC with most other SIVI methods obtaining similar performance levels (with the exception of SM). See Section H.3 for all the implementation details.
Figure 3: Comparison between SIVI methods and MCMC on Bayesian logistic regression problem. (a) shows the marginal and pairwise approximations of posterior of the weights $x_{1},x_{2},x_{3}$, and (b) shows the scatter plot of the correlation coefficient of MCMC ($y$-axis) vs PVI ($x$-axis).